# Single-Image Shadow Removal Using Deep Learning: A Comprehensive Survey

### v1

@@ V-C Computational Complexity
SP+M-Net [51] 0.023 141.18 61.52 PyTorch
DHAN [48] 0.283 21.75 262.87 TensorFlow
EMDN [29] 0.015 10.06 56.29 PyTorch
BMNet [54] 0.036 0.37 10.99 PyTorch
SL ShadowFormer [55] 0.031 11.35 64.60 PyTorch
Inpaint4Shadow [74] 0.006 14.98 81.18 PyTorch
ShadowDiffusion [30] 0.723 60.74 937.15 PyTorch
HomoFormer [78] 0.042 17.81 35.63 PyTorch
UL DC-ShadowNet [58] 0.008 10.59 52.55 PyTorch
ZSL G2R-ShadowNet [63] 0.010 22.76 113.93 PyTorch
BCDiff [64] 31.446 276.41 139218.75 PyTorch
VI Shadow Removal Applications
@@ VI-A Shadow Generation
(c) SP+M-Net [51]
(d) EMDN [29]
(e) G2R-ShadowNet [63]
(f) Inpaint4Shadow [74]
(g) ShadowFormer [55]
(h) ShadowDiffusion [30]
(i) HomoFormer [78]
@@
(c) SP+M-Net [51]
(d) EMDN [29]
(e) G2R-ShadowNet [63]
(f) Inpaint4Shadow [74]
(g) ShadowFormer [55]
(h) ShadowDiffusion [30]
(i) HomoFormer [78]
@@
VI-B Shadow-related Attacks
@@
VII Future Research Directions
As observed from the experimental results presented in Section V, shadow removal remains a challenging research topic, particularly concerning generalization ability. There remains considerable potential for improvement. This section suggests potential future research directions as follows:
VII-A Generalized Shadow Removal
@@
VII-B Interactive Shadow Removal
Interactive shadow removal involves users actively participating in removing shadows from images through a user-friendly interface, typically utilizing tools or algorithms enabling real-time adjustments and feedback. This method grants users enhanced control and customization over the shadow removal process, allowing them to tailor results to their specific preferences and requirements.
@@ VII-B Interactive Shadow Removal
VII-C More Comprehensive Benchmark
@@
VII-D Non-Reference Evaluation Metric
@@
VII-E Extension to Video Shadow Removal
@@
VII-F Large-Scale Training Dataset
While numerous training datasets are available for shadow removal, their scale and diversity in representing various shadow categories do not fully capture the complexity of real shadow degradations, as indicated in Table II. Current deep shadow removal models, as discussed in Section V, encounter challenges in achieving satisfactory performance when confronted with shadow images captured in real-world scenarios. Moreover, the resolution of existing datasets is typically low, such as $480\times 640$, whereas real-world scenarios often involve high-resolution imagery, such as $2k$ or $4k$. Models trained on low-resolution data often struggle to generalize directly to high-resolution scenarios. Therefore, additional efforts are needed to explore the acquisition of large-scale, high-resolution, and diverse real-world paired shadow removal training datasets.
VIII Conclusion

### v2

@@ 5.3 Computational Complexity
6 Shadow Removal Applications
@@ 6.1 Shadow Generation
(c) SP+M-Net
(d) EMDN
(e) G2R-ShadowNet
(f) Inpaint4Shadow
(g) ShadowFormer
(h) ShadowDiffusion
(i) HomoFormer
@@
(c) SP+M-Net
(d) EMDN
(e) G2R-ShadowNet
(f) Inpaint4Shadow
(g) ShadowFormer
(h) ShadowDiffusion
(i) HomoFormer
@@
6.2 Shadow-related Attacks
@@
7 Future Research Directions
As observed from the experimental results presented in Section 5, shadow removal remains a challenging research topic, particularly concerning generalization ability. There remains considerable potential for improvement. This section suggests potential future research directions as follows:
7.1 Generalized Shadow Removal
@@
7.2 Interactive Shadow Removal
Interactive shadow removal [19, 20] engages users in the shadow removal process through a user-friendly interface that allows real-time adjustments and feedback. This approach provides users with greater control and customization, enabling them to tailor the results to their specific preferences and needs. Although deep learning-based methods have shown promise, their application in this interactive domain remains less explored.
@@ 7.2 Interactive Shadow Removal
7.3 More Comprehensive Benchmark
@@
7.4 Non-Reference Evaluation Metric
@@
7.5 Extension to Video Shadow Removal
@@
7.6 Large-Scale Training Dataset
While numerous training datasets are available for shadow removal, their scale and diversity in representing various shadow categories do not fully capture the complexity of real shadow degradations, as indicated in Table 2. Current deep shadow removal models, as discussed in Section 5, encounter challenges in achieving satisfactory performance when confronted with shadow images captured in real-world scenarios. Moreover, the resolution of existing datasets is typically low, such as $480\times 640$, whereas real-world scenarios often involve high-resolution imagery, such as $2k$ or $4k$. Models trained on low-resolution data often struggle to generalize directly to high-resolution scenarios. Therefore, additional efforts are needed to explore the acquisition of large-scale, high-resolution, and diverse real-world paired shadow removal training datasets.
8 Conclusion
@@
Declarations
Conflict of interest. The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Availability of data and materials. This work does not propose a new dataset. All the datasets we used are publicly available.
Code availability. The related benchmark code and repository of this work is released at https://github.com/GuoLanqing/Awesome-Shadow-Removal.